OpenStudy (anonymous):

i really need help pleas Which of the following statements best describes the graph of -5x + 2y = 1? It is a curve joining the points (-5, 2), (2, 3), and (4, 1). It is a curve joining the points (-1, -3), (-1, -3), and (1, 5). It is a straight line joining the points (1, 3), (3, 8), and (-3, -7). It is a straight line joining the points (4, -3), (-1, 2), and (-4, 5).

OpenStudy (anonymous):

You can convert this equation to the slope intercept form. The original equation is in the standard form of Ax + Bx = C

OpenStudy (anonymous):

im sorry that makes no since to me

OpenStudy (anonymous):

slope intercept form is y = mx + b; m is the slope, b is the y-intercept

OpenStudy (anonymous):

-5x + 2y = 1; do you know how to solve for y?

OpenStudy (anonymous):

i dont understand slopes no

OpenStudy (anonymous):

\[-5x + 2y = 1\] \[2y = 5x + 1 \]

OpenStudy (anonymous):

\[y = \frac{ 5 }{ 2 }x + \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

You can rule out the first two options because any equation in the slope intercept form is a linear equation, meaning that there is a straight line connecting the coordinates

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

then

OpenStudy (anonymous):

To find out the points, just plug them into the equation and see if they match

OpenStudy (anonymous):

how do i do that exactly

OpenStudy (anonymous):

Okay let's start with the third option. The coordinates are (1, 3), (3, 8), and (-3, -7) Our equation is \[y = \frac{ 5 }{ 2 }x + \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

ok wat do we do

OpenStudy (anonymous):

Let's start out with the first coordinates (1, 3). Here 1 is the x-value and 3 is the y-value. Plug them into or equation. Is \[\frac{ 5 }{ 2 }(1) + \frac{ 1 }{ 2 } \] equal to our corresponding y-value, 3?

OpenStudy (anonymous):

5/2 means 2.5 right? When you multiply 1 by 2.5, you get 2.5. Add 1/2 which is 0.5, and you get 3. SO far the first coordinates match. Let's get to the other two coordinates of the third option.

OpenStudy (anonymous):

(3, 8); x = 3, y = 8 Is 2.5(3) + 0.5 equal to 8. Yes they do. \[8 = \frac{ 5 }{ 2 }(3) + \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

(-3, -7); x = -3, y = -7 Is 2.5(-3) + 0.5 equal to -7. Yes they do. \[-7 = \frac{ 5 }{ 2 }(-3)+ \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

SO BECAUSE THE VALUES ALL MATCH WITH EACH OTHER IN OUR EQUATION, THE ANSWER IS THE THIRD OPTION

OpenStudy (anonymous):

oh thx

OpenStudy (anonymous):

np